i need help doing these please.
These are suppose to be open sentences.
[3b-5]=7
[w+5]>1

Answer by stanbon(75887)

[3b-5]=7
3b-5=7 or 3b-5=-7
3b=12 or 3b=-2
b=4 or b=-2/3
=========================
[w+5]>1
w+5>1 or x+5<-1
x>-4 or x<-6
------------
Cheers,
Stan H.